I have the hardest time setting up word problems! I can solve them without setting them up but I must show my work.
The weight of a liquid varies directly as its volume V. If the weight of the liquid in a cubical container 4cm on a side is 192g find the weight of the liquid in a cubical container 3cm on a side.

Answer by checkley77(12844)

4^3/192=3^3/x
64/192=27/x
64x=192*27
64x=5,184
x=5,184/64
x=81 is the weight of the liquid in the 3 cm cube.